OpenStudy (anonymous):

what is (5151)+15312

OpenStudy (anonymous):

help

OpenStudy (anonymous):

ergh, seriously? just add both of the data together. 20463

OpenStudy (anonymous):

are you kidding

OpenStudy (anonymous):

no i am not kiding i am 7 years old

OpenStudy (anonymous):

do you have a calculator function on your computer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hello?????

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

okay then, well do you know how to do basic math? 12344 +3433 ________ 15777 You will do basic addition 4+3=7 for the fifth column. 4+3=7 for the fourth column 3+4=7 for the third 2+3=5 for the second 1+0=1 for the first so this question that i am answering would be 15777 just a little longer to add out.

hero (hero):

You are a kid, so you are kidding, lol

OpenStudy (anonymous):

no

hero (hero):

Have you figured this out yet?

OpenStudy (anonymous):

yes why how old are you

hero (hero):

Older than you

hero (hero):

What answer did you get?

OpenStudy (anonymous):

15777

OpenStudy (anonymous):

woah now, that was a 'hypothetical' question i answered to show you how adding large numbers worked.

hero (hero):

lol

OpenStudy (anonymous):

what

hero (hero):

What she means to say is, 15777 is not the answer to your math problem tanya

OpenStudy (anonymous):

oh i mean 20

OpenStudy (anonymous):

you are mean

hero (hero):

I wasn't trying to be

OpenStudy (anonymous):

i am going to tell my big sister

hero (hero):

Okay

hero (hero):